OpenStudy (anonymous):

1. 8y^2-4y+1 divided by 2y-1 2. 2a^2+a+3 divided by a-1 3. 6x^3+11x^2-4x-4 divided by 3x-2 4. 6x^2+11x-35 divided by 3x-5 5. x^2-4 divided by x-1 6. y^3-4y^2+7y-6 divided by y-2 7. x^3+y^3 divided by x-y 8. a^4+4b^4 divided by a^2-2ab+2b^2 9. a^2^n-a^n-6 divided by a^n+8 thanks guyss

OpenStudy (lgbasallote):

seeing as your problems arent cancelable...ill do this the long method... 1. 8y^2 - 4y + 1/2y - 1 = 4y + 1/(2y-1) 2.2a^2 +a +3/a -1 = 2a + 3 + 6/(a-1) those are the first two...please wait for the others

OpenStudy (lgbasallote):

3. 6x^3 + 11x^2 - 4x - 4/3x -2 = 2x^2 + 5x + 2